OpenStudy (anonymous):

How can you write the expression with rationalized denominator? sqrt 3 - sqrt 6/ sqrt 3 + sqrt 6

OpenStudy (agent0smith):

You still are not using parentheses... :P

OpenStudy (anonymous):

Is doesnt have parenthesis :P

OpenStudy (agent0smith):

It doesn't need them, since it's probably written as a fraction. But when you're writing them, you need to show them. Anytime you have things in the numerator and denominator, there needs to be parentheses around the entire numerator, and entire denominator.

OpenStudy (anonymous):

Okay cx lol so ...how to i get my answer? cx

OpenStudy (agent0smith):

if it's this,\[\Large \frac{\sqrt 3 - \sqrt 6} { \sqrt 3 + \sqrt 6 }\]then you need parentheses, like so: (sqrt 3 - sqrt 6) / (sqrt 3 + sqrt 6)

OpenStudy (agent0smith):

you'll have to multiply by the conjugate of the denominator like so\[\Large \frac{\sqrt 3 - \sqrt 6} { \sqrt 3 + \sqrt 6 }*\frac{ \sqrt 3 - \sqrt 6 }{ \sqrt 3 - \sqrt 6 }\]

OpenStudy (anonymous):

How do i multiply that?

OpenStudy (agent0smith):

use FOIL, ie expand it out\[\Large \frac{(\sqrt 3 - \sqrt 6)} { (\sqrt 3 + \sqrt 6) }*\frac{ (\sqrt 3 - \sqrt 6) }{ (\sqrt 3 - \sqrt 6) } = \]

OpenStudy (agent0smith):

this is not going to be easy to follow, sorry. You'll have to try it yourself on paper \[= \Large \frac{\sqrt 3*\sqrt 3 - \sqrt 6 \sqrt 3 - \sqrt 6 \sqrt 3+\sqrt 6 \sqrt 6} { \sqrt 3 \sqrt 3+ \sqrt 3\sqrt 6 - \sqrt 6 \sqrt 3 +\sqrt 6 \sqrt 6 }\]

OpenStudy (agent0smith):

\[= \Large \frac{3 -2 \sqrt {18}+6} { 3+ 6 }\]

OpenStudy (agent0smith):

All i did was FOIL it, multiply it all together. Now just gotta simplify

OpenStudy (anonymous):

sqrt(3) - sqrt(6) ------------- sqrt(3) + sqrt(6) -1 - 2 sqrt(18) ------------- 3 -3 - 2 sqrt(18) ------------- 9 -3 + 2 sqrt(2) <---- answer 9 - 2 sqrt(18)

OpenStudy (anonymous):

thanks :P

OpenStudy (anonymous):

aha np :)